Project

General

Profile

HallMeeting-2014-02-12

1. Muon Beamline – Henry Nebrensky
*Decay solenoid:*RGA is cleaned ready for use
Problem with one flow meter – Mike contact DL, but can start cooldown
Aim to complete powering tests before end-March
3 A/C units working
Repaired current lead has passed insulation test

Conventional magnets:
Henry – chase up handle on Q8 PSU – stands

Target: Send irradiated fibres to Sheffield – Andy - stands
Request to put R78 target onto MICEnet, but be wary of security – Pierrick investigate other solutions - stands
Offline target inspection takes place 19th Feb – Target Group
Also run online target with frame raised on 19th – raise with MEMO - Henry

Proton absorber:
Installation plan is made, request for survey to be made through Tim Hayler.

2. FC, liquid hydrogen and absorber integration - Steve Watson

Absorber pipes have been made to fit, still need to fix absorber level sensor connector – Steve/DL
Rootes blowers will need high volume N2 supply in R5.2 – revive CryoEase idea? Andy

3. Tracker activity in MICE Hall and R1/R9 – Craig Macwaters

Just fitted tracker model in OK, John Palin to measure next week, then build platforms, etc. and refit helium window.
Schedule going fine for April 4th
Suggest joining FC meeting as required when parallel working in R9 starts
Downstream tracker schedule has been compiled
Length decision has been made for spare lightguides; should someone place order?
Light for darkroom needed for March 10th – Simon, but watch effort clash with RF PSU – use Estates Helpdesk if so

4. Installation of Step IV – Andy Nichols.

West mezzanine completion – 1st week March
Jason – chase start of South ramp cutting – meet Kier tomorrow, Quote next week
South mezzanine mod drawings late – Jason chase OPUS directly – in progress
Jason – request trench cable repositioning through ISIS Footprints – stands, also find spec of ISIS control cables and susceptibility to magnet quench – Field Mitigation Group – important action! - Fit brass interlayer pads to downstream SS support frame – Mark S – low priority
Add Craig to joint infrastructure meeting - Andy
Remove tracker readout cryostat from R5.2 ASAP - Mark

5. Hall Infrastructure – Andy Nichols

RR2 cable channels being widened today
Meeting with Mr Floor this week to route raised floor supports round cable trays in corridor for RR2
Transfer responsibility for air supply from Tim Hayler – Andy
Neutron monitor now working on mains supply – return portable ones to Frances Burge

6. Alignment and survey – Tim Hayler

Measure tracker model in situ – Tim make request
Proton absorber survey - Steve

7. Electrical Infrastructure – Andy Gallagher

Green text: action complete, delete from list
Blue text: waiting for information/parts
Red text: normal work in progress
Install power feeds to RF rack positions. Cables ordered. Drawing issued. Mark up Rack positions and have floor plates cut then install supply cables. On Hold for STEP IV installation

• Re-position 208v Transformer further east in the trench. Re-position into RR2.
• Re-position D13 on North mezz extn. Waiting on Craig’s rack position changes.
Install network cables & limit cable for Proton Absorber.
Remove webcam from west wall when access is available from scaffold.
R9 Dark Room Lights?
Install air lines and limits cable for Proton Absorber.
Install cable tray to RR2.
Install distribution board in RR2 when access is available.
Install cable tray in RR2 when steel plinths are in position.
EMR rack – install filtered air inlet and reverse fans.

Add power to webcam on North mezzanine to list
Mobile phone repeater needs fixing

8. MLCR, controls and networking, Online – Craig Macwaters/Henry Nebrensky

More keys required for MLCR – Andy – stands
MLCR A/C supply confirmed not on substation 25 supply – chase with Gordon Stanford – Craig – has written to him
Change A/C filters before end FY - John
VME crates – hoovered out, two working, one failed, Craig getting quote for new equipment miceioc6 crate has replaced miceioc1 – trying to reduce number of crates - Yordan
Looking at routing structured network cabling – Craig/Simon

9. PPS –- Craig Macwaters
Potterax door broken John /Craig

10. RF progress at DL and RAL – Tim Stanley

Nothing to report

11. Detector preparations and running

Nothing to report

12. R9 activity – Steve Watson

Both Rootes blowers tested OK
FC#1 at 70K, internal leak no worse
Fuse box clearance done
Designate R9 bottle delivery area - Andy
aig Macwaters Potterax door bolts broken – John/Craig – Estates helpdesk?

13. Effective dates of next shutdown and job list –All.
4th April – 6th May 2014, (maintenance days: 26th February, 19th March
Synchrotron phone link Neutron monitor calibration Proton absorber Drip trays for quads/dipole Clean up DSA Target chiller – talk to Paul Hodgson
14. A.O.B
None
15. Date of next meeting
Wednesday, February 19th, G-06, R66, 13:30 local time++.

Hall-meeting12-feb.pdf - Notes as PDF (240 KB) Govans, John, 13 March 2014 16:51